Please help me solve this equation :One pipe can fill a tank in 5 hours less than another. Together they
can fill the tank in 5 hours. How long would it take each alone to fill the tank ? Compute the answer
to two decimal places.

Found 2 solutions by ikleyn, jorel1380:
Answer by ikleyn(52787)

.
Solved in
https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1074980.html

https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1074980.html



Answer by jorel1380(3719)

Let n be the time it takes for the first pipe. Then the second pipe would be n+5. So:
1/n + 1/n+5=1/5
5/n + 5/n+5=1
5n+25+5n=nē+5n
nē-5n-25=0
Using the quadratic formula, we get one positive root of 8.09016994375. So the first pipe takes 8.01 hours, while the second pipe takes 13.01 hours. ☺☺☺☺